OpenStudy (anonymous):

Help? Find all solutions to the equation in the interval [0, 2π). cos 2x - cos x = 0

myininaya (myininaya):

So we could express this in a quadratic like form. Use the identity: \[**\cos(2x)=2\cos^2(x)-1 \] This is how I got it: \[\cos(2x)=\cos^2(x)-\sin^2(x)=\cos^2(x)-(1-\cos^2(x))=2\cos^2(x)-1 \]

myininaya (myininaya):

So you are really asked to solve \[2\cos^2(x)-\cos(x)-1=0\] And you should know how to solve \[2u^2-u-1=0\]

OpenStudy (anonymous):

So I factor it? @myininaya

myininaya (myininaya):

Yep or quadratic formula.

OpenStudy (anonymous):

Awesome.

OpenStudy (anonymous):

Hey... I got -1/2, and 1, but how do I get the answers like my choices? They have pi and such in them.. I don't get it..

myininaya (myininaya):

ok so you solve for cos(x)

myininaya (myininaya):

now you need to solve for x

myininaya (myininaya):

\[\cos(x)=\frac{-1}{2} \text{ or } \cos(x)=1 \]

myininaya (myininaya):

Solve both of those equations

OpenStudy (mathmale):

Hi, BatGirl: If you have a TI-84 calculator, you could type in -1 cos 1, which would result in your obtaining the angle x = 0. Or you could graph y=cos x on the given interval, and note from that x=0 is the only x value on that interval at which cos x = 1.

OpenStudy (anonymous):

Wait, I could type in what? @mathmale

myininaya (myininaya):

Or you could use the unit circle.

myininaya (myininaya):

http://en.wikipedia.org/wiki/Unit_circle

OpenStudy (anonymous):

\[A) 0, 2\pi/3, 4\pi/3\] \[B) \pi/6, 5\pi/6, 3\pi/2\] \[C) 0, \pi/2, 7\pi/6, 11\pi/6\] D) No solution

myininaya (myininaya):

For what angles on your unit circle is the cosine value 1 and -1/2?

myininaya (myininaya):

Just like the x coordinates on your table/unit circle

OpenStudy (anonymous):

pi/3, and pi/2?

myininaya (myininaya):

at pi/2 cos is 0 at pi/3 cos is 1/2

OpenStudy (anonymous):

Ugh. I'm not very good with this circle thing..

myininaya (myininaya):

I'm asking you to look at that circle and tell me for what angles is the x-coordinate -1/2 or 1

OpenStudy (anonymous):

So it's 0, and..

OpenStudy (anonymous):

5pi/3?

myininaya (myininaya):

Look to see where you have the x-coordinate is 1 first? Can you name these values? yes at 0 we have cos is 1 now also when is the x-coordinate -1/2?

myininaya (myininaya):

Look at the x-coordinates until you see -1/2 you should see and angle near by corresponding to that point I'm asking you for that angle

myininaya (myininaya):

There is actually two angles in between 0 and 2pi so that cos value is -1/2

OpenStudy (anonymous):

300?

myininaya (myininaya):

that is 1/2 not -1/2

myininaya (myininaya):

So looking at that circle you don't see any x-coordinates being -1/2? Do you know which one is the x-coordinate? The very first value of any rectangular point is the x-value as long as we are treating the horizontal axis as the x-axis which is what we normally do.

OpenStudy (anonymous):

Oh! Oops! 240.

myininaya (myininaya):

which 4pi/3 in radians now one more angle will give you the cos is -1/2

myininaya (myininaya):

hint look in the second quadrant

OpenStudy (anonymous):

120, or 2pi/3. But there's another one.. How do we know to only use those two?

myininaya (myininaya):

So we know to use 0, 4pi/3, and 2pi/3 because those are the only angles between 0 and 2pi that give us cos is either 1 or -1/2

OpenStudy (anonymous):

Wait, nevermind, that one is a y-value. My bad.

OpenStudy (anonymous):

What does this mean? Rewrite with only sin x and cos x. cos 2x + sin x

myininaya (myininaya):

Well your expression is in terms of cos(2x) and sin(x) not sin(x) and cos(x)

myininaya (myininaya):

You have to use an identity to rewrite cos(2x)

myininaya (myininaya):

the same identity I gave you above.

OpenStudy (anonymous):

That's what I thought, but none of my answers have a cos in them... :c

myininaya (myininaya):

Well rewrite it using sine then

myininaya (myininaya):

you know cos^2(x)=1-sin^2(x)

OpenStudy (anonymous):

It's not squared, just has a 2x though.

myininaya (myininaya):

? we said earlier cos(2x)=2cos^2(x)-1 I'm saying use this identity cos^2(x)=1-sin^2(x) to rewrite cos(2x) in terms of sin

OpenStudy (anonymous):

Oh.. My bad again. I think I can get that. Thanks so much! God bless you!

OpenStudy (mathmale):

Hi, BatGirl: If you have a TI-84 calculator, you could type in -1 cos 1, which would result in your obtaining the angle x = 0. By that I meant the inverse cosine of 1. -1 If you are told that the cosine of a certain angle is 1, then cos 1 on your calculator will produce the angle in either degrees or radians, depending on the MODE you've selected.